I can't figure out how to solve this problem:6(x+3)=9x
Answer by checkley77(12569)

6(x+3)=9x
6x+18=9x
6x-9x=-18
-3x=-18
x=-18/-3
x=6 ans.
Proof:
6(6+3)=9*6
6*9=54
54=54